[figure: 9.7 cm]

STOICHIOMETRY AND MAGNETIC AND STRUCTURAL PROPERTIES OF PURE BISMUTH IRON GARNET

T. Okuda, N. Adachi, and T. Kaneuji Nagoya Institute of Technology, Nagoya, 466-8555 Japan

Pure Bi_{3}F e_{5}O_{12} (BIG) is only obtainable as a thin film by direct vapor-phase epitaxial growth during deposition. As the crystal growth is taken place on a garnet substrate kept at around 500^{∘}C in a reduced atmosphere and Bi is highly volatile, the stoichiometric composition of BIG (Bi/F e = 0.6) may be hardly assured. In addition, it is very difficult to distinguish the compositional ratio of BIG phase from that of the whole film. Therefore we tried to evaluate the quality of BIG phase by comparing the physical properties of the films grown using sputtering targets with various Bi/Fe ratio. Among the films in which only BIG phase is observed by XRD means, the film with Bi/F e = 0.53 is revealed to have the best quality from the following point of view; the narrowest FWHM of X-ray rocking curve and FMR line width, and the largest magnetization and Faraday rotation angle. This result suggests that BIG phase contains a considerable amount of vacancy at Bi sites.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : oral

Corresponding author : T. Okuda

Address for correspondence : Nagoya Institute of Technology Gokiso-cho, Showa-ku, Nagoya 466-8555 Japan

Email address : okudatks@mse.nitech.ac.jp